A car moving at 30 miles/hour has roughly 100,000 J of kinetic energy. You need to keep your foot on the gas to maintain your speed (assume a flat road). The gas you burn represents 500 J of chemical energy every second. (show work)


a. After 5 minutes, how much energy have you generated by burning gasoline?
b. How much kinetic energy does your car have at this time?
c. If this number is not 100,000 J, explain why. If this number is 100,000 J, explain what happened to the extra energy you calculated in part a.
Answer #1

a. Energy burnt per second=500 J.

So, energy burnt in 5 minutes=500*5*60=150,000 J

b)According to the problem,the car maintains its speed.

kinetic energy=1/2mv2 where m is mass and v is velocity. v is constant, as the car maintains its speed , and mass of car, m , is obviously constant.

So,kinetic energy is constant. Hence,kinetic energy=100,000 J

c)Extra energy supplied in part a is used to overcome friction and is lost as heat. So, it does not result in increase in kinetic energy.